a 17.92 g sample of copper was submerged in water, and the volume reading rose by 2.00 mL. what is the density of the copper sample?


1
Expert's answer
2021-09-28T07:38:19-0400

Solution:

Density = Mass / Volume = g / cm3


Since the volume of water rose by 2.00 mL, then the volume of copper is 2.00 mL.


Convert mL to cm3:

1 mL = 1 cm3

Hence,

(2.00 mL) × (1 cm3 / 1 mL) = 2.00 cm3


Density = (17.92 g) / (2.00 cm3) = 8.96 g/cm3

Density = 8.96 g/cm3


Answer: The density of the copper sample is 8.96 g/cm3.